# A Practical Review of Mechanistic Interpretability for Transformer-Based Language Models

### v2

@@ 4.2.3 Technical Advancements
Intervention-based methods are widely used to causally validate hypotheses about how specific LM components contribute to its behavior. For instance, Causal Scrubbing Chan et al. (2022a) is a popular technique for formalizing and validating a hypothesis about the function of an LM component for the given behaviors. Specifically, causal scrubbing involves formalizing a hypothesis $(G,I,c)$, where $G$ is the model’s original computational graph, $I$ is an interpretation graph that reflects the hypothesized roles of the model’s components, and $c$ is a function that maps between nodes of $I$ to $G$. For example, when an LM is tasked to perform a two-digit addition, the interpretation graph $I$ could specify the hypothesized computational pathway, where only certain transformer components are used for implementing this addition while others do not contribute to this task. To validate that $I$ is a faithful interpretation of $G$ for a given behavior, resampling ablations of component activations in $G$ is performed. In our example, it means that replacing activations of model components that do not contribute to the addition task should not change the LM’s prediction behavior. This provides a more rigorous way to check our intuitions about how models work, compared to just looking at the model or manually changing its parts. Similarly, Interchange Intervention Geiger et al. (2021) is another hypothesis verification technique that similarly formulates the hypothesis by representing the model $M$ as a high-level conceptual model and performing resampling ablation on the low-level (i.e., the original) model to verify the hypothesis.
@@
The intervention technique has also been adapted by Pal et al. (2023); Ghandeharioun et al. (2024) to decode features from intermediate activations, which can be viewed as augmented vocabulary projection (Section 4.1) with more powerful expressivity. For instance, to decode features from the end-position activation $h_{n}^{l}$, the fixed prompt causal intervention Pal et al. (2023) runs another LM inference on a fixed generic prompt (e.g., “Tell me something about”) with its activation at the same location being patched by the target one, and then let the LM continue the generation. This generation can then be viewed as the projected features encoded in the activation $h_{n}^{l}$. When this fixed prompt is replaced by one learned for effective information elicitation, it results in the variant of learned prompt causal intervention (or future lens, as the initial goal of Pal et al. 2023 was to probe the future tokens from the end-position activation). Concurrently, Patchscopes Ghandeharioun et al. (2024) provides a unifying framework that similarly combines vocabulary projection with intervention methods but is more generalized than the approaches of Pal et al. (2023). Specifically, Patchscopes defines a source prompt and a source LM, which produces the activation to be interpreted, as well as a target prompt and a target LM, which facilitates the activation interpretation. Patchscopes works by patching the activation derived from the source to the corresponding activation in the target. By carefully designing the target prompt (also termed as “inspection prompt”), Patchscopes allows one to decode the interested feature information. For instance, to check if an LM’s activation obtained from the end position of the source prompt “Amazon’s former CEO” encodes the feature of “Jeff Bezos”, an inspection prompt (or $X_{corrupt}$) can be “cat $\rightarrow$ cat; 135 $\rightarrow$ 135; hello $\rightarrow$ hello; ?”, which comprises few-shot demonstrations encouraging the model to explain its internal representation, with the final “?” serving as a placeholder. If the LM with the activation at the position of “?” patched by the source activation can generate “Jeff Bezos” as the continuation, we can infer that the activation contains information about this person’s name.
@@ 4.3.1 Basic Concepts
Here, the encoder consisting of weights $W_{enc}\in\mathbb{R}^{d\times s}$, a bias term $b_{enc}\in\mathbb{R}^{s}$, and a non-linear activation function $g$ (e.g., ReLU Agarap (2018)), maps the input activation $h\in\mathbb{R}^{d}$ to sparse activation $f(h)\in\mathbb{R}^{s}$. Subsequently, the decoder consisting of weights $W_{dec}\in\mathbb{R}^{s\times d}$ and a bias term $b_{dec}\in\mathbb{R}^{d}$, takes the sparse activation $f(h)$ from the encoder and reconstructs the input activation $\hat{h}$.
@@ 4.3.3 Technical Advancements
SAEs have emerged as a dominant approach for conducting open-ended feature studies (Section 5.1.2) and have garnered significant attention within the MI community. The technical advancements in SAEs can be broadly categorized into two categories: (1) Balancing the trade-off between reconstruction and sparsity loss; (2) Discovering features that are functionally useful for the LM to implement downstream applications
@@
The loss function of SAEs consists of two objectives: reconstruction loss, i.e., the sparse representation should accurately preserve information from input activation, and sparsity loss, i.e., only a small number of elements in the sparse representation should be active for any given input activation. These two objectives can conflict, as greater sparsity can reduce reconstruction loss Gurnee (2024). For instance,  Wright & Sharkey (2024) noted that $L_{1}$ penalty in sparsity loss can lead to feature suppression, a systematic underestimation of feature activation magnitudes, particularly those with weak activations but high frequency, negatively impacting reconstruction loss. To address this, a line of research has focused on improving SAE architectures and training methods to better balance the trade-off between sparsity and reconstruction loss Gao et al. (2024); Rajamanoharan et al. (2024b); Erichson et al. (2019). For instance, Gao et al. (2024) proposed TopK SAE that enforces sparsity by selecting only the K most active features and setting the rest to zero, where the sparse representation is obtained by $f(h)=\texttt{TopK}(g(W_{enc}h+b_{enc}))$. This approach eliminates the need for an $L_{1}$ penalty to achieve sparsity, mitigating the issue of feature suppression. Similarly, Rajamanoharan et al. (2024b) introduced JumpReLU SAEs, which leverage the JumpReLU function Erichson et al. (2019) in place of the ReLU function. The JumpReLU function sets activations below a positive threshold to zero, effectively removing false positives and increasing sparsity, while leaving activations above the threshold intact, thus preventing feature suppression and improving fidelity. Finally, Gated SAEs Rajamanoharan et al. (2024a) address the issue of feature suppression by using a gated ReLU encoder that decouples the detection of active features from the estimation of their magnitudes, and applying the $L_{1}$ penalty only to the feature detection. This decoupling allows the Gated SAE to achieve sparsity without underestimating feature magnitudes.
@@
The primary goal of training SAEs is to discover functionally-important features used by LMs to implement their behavior. Specifically, a feature is considered functionally important if it facilitates the explanation of the model behavior on the training distribution. However, standard SAEs trained to minimize the reconstruction and sparsity loss may not strongly correlate with the goal of discovering functionally important features Braun et al. (2024); Makelov et al. (2024). To address this, Braun et al. (2024) proposed end-to-end (e2e) SAEs, which are trained to minimize the KL divergence between the output distributions of the original model and the model with SAE-generated activations inserted, instead of minimizing the reconstruction loss. This approach ensures that e2e SAEs are optimized to identify features that influence the model’s predictions, rather than focusing on features that only accurately reconstruct activations.
@@ 5 A Beginner’s Roadmap to MI
A key motivation for this survey is to provide a friendly guide for researchers and developers interested in MI to quickly pick up the field. To this end, we provide a beginner’s roadmap in Figure 8, 11, and 13, where we categorize MI research into three categories – feature study, circuit study, and the study of universality, corresponding to the three objects of studies (Section 3.1). Each category is further divided into sub-categories with their corresponding MI workflow and associated techniques.
@@ 5.1 Feature Study
A targeted feature study aims to determine whether a pre-defined feature is encoded in the representation of an LM. These pre-defined features are typically intuitive guesses or hypotheses made by humans, and we perform a targeted feature study when we want to verify the presence or absence of these features. Therefore, a general research question in a targeted feature study is, “Does this LM representation (or a subset of its neurons) encode the feature $F$?”
@@ 5.1.1 General Workflow for Targeted Feature Study
A targeted feature study begins with a hypothesis about the presence or absence of a pre-defined feature $F$ in an LM representation. For example, suppose we are interested to know whether an LM activation encodes the “is_python_code” feature in its activations when the input contains a Python code snippet. Then, our hypothesis would be: “Does this LM activation (e.g., $h_{i}^{1}$) encode the ’is_python_code’ feature?” We can refine our hypothesis to localize the feature encoding to a more granular level: “Does this neuron in this activation (e.g., Neuron 1 in $h_{t}^{1}$) encode “is_python_code” feature?”. Note that we can iteratively pose this question to all the LM activations when we want to identify a complete set of activations that encode the targeted feature.
@@
To verify whether or not a pre-defined feature exists in a given LM representation, there are typically two workflows to consider. The first workflow directly probes the LM representation and inquires the presence of the targeted feature $F$, whereas the second workflow follows a similar procedure as the open-ended feature study, which first identifies all features encoded in the representation and then confirms whether the feature $F$ is among them. Below, we mainly present the first workflow. Details of the second workflow will be discussed in Section 5.1.2. For the first workflow, one option is to use the Probing method (Section 4.4.1) to verify the hypothesis, where a probe determining the existence of the targeted feature $F$ will be trained. Alternatively, PatchScope Ghandeharioun et al. (2024) is a flexible framework integrating vocabulary projection and intervention, and can also be used for verifying the above hypothesis. This method is especially useful to verify whether certain attributes associated with a given subject (e.g., attributes associated with the “largest city of” from the representation of “United States”) are encoded in the LM representation. In comparison to Probing, PatchScope enjoys the benefit of needing no additional training. However, it relies on humans to craft an expressive inspection prompt capable of eliciting the desired feature, which can be particularly challenging for abstract features. Therefore, the choice of these two techniques often depends on the type of feature being investigated and the availability of computing and data resources.
@@ 5.1.2 General Workflow for Open-ended Feature Study
Vocabulary projection methods (Section 4.1), neuron activation visualization (Section 4.4.2), and SAE (Section 4.3) are commonly used techniques for generating intermediate explanations. They are referred to as “intermediate explanations” because they require further interpretation during the Explain step to uncover features.
@@
Neuron activation visualization (Sec 4.4.2) is another technique that involves developing an interface to visually highlight the text or tokens which elicit activity of a specific neuron, to create an intermediate explanation, as illustrated in Figure 9. The highlighted text snippets are typically sampled from a large corpus, focusing on instances where the neuron exhibits the highest activation. The idea here is that, for example, if a neuron consistently activates for text written in French, one can then label the neuron as encoding a French text detection feature. It is important to note that neuron activation visualization is based on an important assumption that features are represented by neurons in a one-to-one correspondence. However, subsequent studies Elhage et al. (2022a); Marks et al. (2024b) have found polysemantic neurons, i.e., neurons that activate for multiple unrelated features, rendering the assumption of one-to-one correspondence to be false. The discovery of polysemantic neurons also complicates the Explain step, since polysemantic neurons can easily confuse the explainer when they activate for multiple unrelated features simultaneously.
To address this challenge, subsequent studies have proposed training SAEs (Section 4.3) that will transform the $n$-dimensional original activation into a higher $k$-dimensional sparse activation. The goal of this projection is to generate a new representation that faithfully retains all the features from the original activation but encodes them in a way that each neuron corresponds to a single feature, resulting in a representation composed of monosemantic neurons. The neuron activation visualization can then be applied to the sparse activation, rather than the original activation, for generating more interpretable intermediate explanations. Research on SAEs is still in its early stages, but it has shown promising results by discovering features that were not identified through neuron activation visualizations Bricken et al. (2023); Templeton et al. (2024), which we will summarize in Section 7.1.
Figure 9: The neuron activation visualization developed by Elhage et al. (2022a) for labeling neurons with a feature by analyzing the text on which they activate.
Although each of these techniques – vocabulary projection methods, neuron activation visualization, or SAEs, can be used individually to generate the intermediate explanation in the Observe step, in practice they are often combined to provide a more complete explanation of the LM activation. A user interface incorporating intermediate explanations from multiple techniques is shown in Figure 10. This may also include other explanations, such as activation density, which indicates how sparsely a neuron or feature activates for a given dataset. The availability of all these explanations can help the explainer more confidently identify and explain the features encoded in the corresponding activations.

### v4

@@ 5.2.3 Technical Advancements
Intervention-based methods are widely used to causally validate hypotheses about how specific LM components contribute to its behavior. For instance, Causal Scrubbing Chan et al. (2022a) is a popular technique for formalizing and validating a hypothesis about the function of an LM component for the given behaviors. Specifically, causal scrubbing involves formalizing a hypothesis $(G,I,c)$, where $G$ is the model’s original computational graph, $I$ is an interpretation graph that reflects the hypothesized roles of the model’s components, and $c$ is a function that maps between nodes of $I$ to $G$. For example, when an LM is tasked to perform a two-digit addition, the interpretation graph $I$ could specify the hypothesized computational pathway, where only certain transformer components are used for implementing this addition while others do not contribute to this task. To validate that $I$ is a faithful interpretation of $G$ for a given behavior, resampling ablations of component activations in $G$ are performed. In our example, it means that replacing activations of model components that do not contribute to the addition task should not change the LM’s prediction behavior. This provides a more rigorous way to check our intuitions about how models work, compared to just looking at the model or manually changing its parts. Similarly, Interchange Intervention Geiger et al. (2021) is another hypothesis verification technique that similarly formulates the hypothesis by representing the model $M$ as a high-level conceptual model and performing resampling ablation on the low-level (i.e., the original) model to verify the hypothesis.
@@
The intervention technique has also been adapted by Pal et al. (2023); Ghandeharioun et al. (2024) to decode features from intermediate activations, which can be viewed as augmented vocabulary projection (Section 5.1) with more powerful expressivity. For instance, to decode features from the end-position activation $h_{n}^{l}$, the fixed prompt causal intervention Pal et al. (2023) runs another LM inference on a fixed generic prompt (e.g., “Tell me something about”) with its activation at the same location being patched by the target one, and then let the LM continue the generation. This generation can then be viewed as the projected features encoded in the activation $h_{n}^{l}$. When this fixed prompt is replaced by one learned for effective information elicitation, it results in the variant of learned prompt causal intervention (or future lens, as the initial goal of Pal et al. 2023 was to probe the future tokens from the end-position activation). Concurrently, Patchscope Ghandeharioun et al. (2024) provides a unifying framework that similarly combines vocabulary projection with intervention methods but is more generalized than the approaches of Pal et al. (2023). Specifically, Patchscope defines a source prompt and a source LM, which produces the activation to be interpreted, as well as a target prompt and a target LM, which facilitates the activation interpretation. Patchscope works by patching the activation derived from the source to the corresponding activation in the target. By carefully designing the target prompt (also termed as “inspection prompt”), Patchscope allows one to decode the interested feature information. For instance, to check if an LM’s activation obtained from the end position of the source prompt “Amazon’s former CEO” encodes the feature of “Jeff Bezos”, an inspection prompt (or $X_{corrupt}$) can be “cat $\rightarrow$ cat; 135 $\rightarrow$ 135; hello $\rightarrow$ hello; ?”, which comprises few-shot demonstrations encouraging the model to explain its internal representation, with the final “?” served as a placeholder. If the LM with the activation at the position of “?” patched by the source activation can generate “Jeff Bezos” as the continuation, we can infer that the activation contains information about this person’s name.
@@ 5.3.1 Basic Concepts
Here, the encoder consisting of weights $W_{enc}\in\mathbb{R}^{d\times s}$, a bias term $b_{enc}\in\mathbb{R}^{s}$, and a non-linear activation function $g$ (e.g., ReLU Agarap (2018)), maps the input activation $h\in\mathbb{R}^{d}$ to sparse activation $f(h)\in\mathbb{R}^{s}$. The size of the sparse representation $s$ is a hyperparameter, decided by the researcher. Subsequently, the decoder consisting of weights $W_{dec}\in\mathbb{R}^{s\times d}$ and a bias term $b_{dec}\in\mathbb{R}^{d}$, takes the sparse activation $f(h)$ from the encoder and reconstructs the input activation $\hat{h}$. The decoder is also commonly referred to as the dictionary, as it comprises a set of learned feature vectors encoded in its weight matrix after training. Consequently, the size of the sparse representation $s$ is referred to as the dictionary size.
@@ 5.3.2 Technique-specific Interpretation
Approach Basic Concepts Technical Advancements
Standard SAE Bricken et al. (2023) Train a standard SAE with ReLU activation and $L1$ sparsity that maps model activations into higher-dim sparse activations with more monosemantic units, which helps open-ended feature discovery. N/A
TopK SAE Gao et al. (2024) Enforce sparsity without using L1 penalties by retaining only the top-K activations per input, reducing feature suppression caused by L1 regularization. Improve sparsity reconstruction trade-off.
JumpReLU SAE Rajamanoharan et al. (2024b) Replace ReLU with the JumpReLU activation function, which sets an activation threshold to remove false-positive sparse activation. Improve sparsity reconstruction trade-off.
Gated SAE Rajamanoharan et al. (2024a) Introduce a modified encoder architecture with two separate pathways: a gating path, which determines active features and is subject to a sparsity penalty, and a magnitude path, which estimates the strength of each active feature and is not penalized for sparsity Improve sparsity reconstruction trade-off.
Matryoshka SAE (Bussmann et al., 2025) Trains multiple nested sub-SAEs of increasing dictionary sizes simultaneously to learn a common feature space that consists of both general and specific features. Improve sparsity reconstruction trade-off.
End-to-End SAE Braun et al. (2024) Minimize KL divergence between the original model’s output and that of the model with SAE-generated activations, rather than reconstruction loss. Improve discovery of functionally important features.
Table 6: Representative approaches in the technical advancements of sparse autoencoders (SAE).
@@ 5.3.3 Technical Advancements
SAEs have emerged as a dominant approach for conducting open-ended feature studies (Section 6.1.2) and have garnered significant attention within the MI community. The technical advancements in SAEs can be broadly categorized into two categories: (1) Balancing the trade-off between reconstruction and sparsity loss; (2) Discovering features that are functionally useful for the LM to implement downstream applications. We summarize representative approaches discussed in the technical advancements of SAEs in in Table 6.
@@
The loss function of SAEs consists of two objectives: reconstruction loss, i.e., the sparse representation should accurately preserve information from input activation, and sparsity loss, i.e., only a small number of elements in the sparse representation should be active for any given input activation. These two objectives can conflict, as greater sparsity can reduce reconstruction loss Gurnee (2024). For instance,  Wright & Sharkey (2024) noted that $L_{1}$ penalty in sparsity loss can lead to feature suppression, a systematic underestimation of feature activation magnitudes, particularly those with weak activations but high frequency, negatively impacting reconstruction loss. To address this, a line of research has focused on improving SAE architectures and training methods to better balance the trade-off between sparsity and reconstruction loss Gao et al. (2024); Rajamanoharan et al. (2024b); Erichson et al. (2019). For instance, Gao et al. (2024) proposed TopK SAE that enforces sparsity by selecting only the K most active features and setting the rest to zero, where the sparse representation is obtained by $f(h)=\texttt{TopK}(g(W_{enc}h+b_{enc}))$. This approach eliminates the need for an $L_{1}$ penalty to achieve sparsity, mitigating the issue of feature suppression. Similarly, Rajamanoharan et al. (2024b) introduced JumpReLU SAEs, which leverage the JumpReLU function Erichson et al. (2019) in place of the ReLU function. The JumpReLU function sets activations below a positive threshold to zero, effectively removing false positives and increasing sparsity, while leaving activations above the threshold intact, thus preventing feature suppression and improving fidelity. Gated SAEs Rajamanoharan et al. (2024a) address the issue of feature suppression by using a gated ReLU encoder that decouples the detection of active features from the estimation of their magnitudes, and applying the $L_{1}$ penalty only to the feature detection. This decoupling allows the Gated SAE to achieve sparsity without underestimating feature magnitudes. In addition to feature suppression, the other known issues include (1) feature splitting (Bricken et al., 2023), where a general features (e.g., punctuation marks) fragments into more specific features (comma, period, question mark); (2) feature absorption (Chanin et al., 2024), where parent feature only partially splits leading to general feature with holes (e.g., a feature that activates on all tokens starting with an “E”, except if the word is “Elephant”); and (3) feature composition (Anders et al., 2024), where distinct features are entangled into a single feature (e.g., “red triangle” instead of “red” and “triangle” separately). To address these issues, Bussmann et al. (2025) proposed Matryoshka SAE, a novel hierarchical approach to SAE training that mirrors the hierarchical structure of real-world features. Specifically, Matryoshka SAE trains multiple nested sub-SAEs of increasing dictionary sizes simultaneously to learn a common feature space that consists of both general and specific features. Each sub-SAE is optimized to reconstruct the input using only a subset of the total latents. This design prevents later, more specialized features from absorbing the roles of earlier, more general ones, thereby regularizing the SAE to capture features across multiple levels of abstraction.
@@
The primary goal of training SAEs is to discover functionally-important features used by LMs to implement their behavior. Specifically, a feature is considered functionally important if it facilitates the explanation of the model behavior on the training distribution However, standard SAEs trained to minimize the reconstruction and sparsity loss may not strongly correlate with the goal of discovering functionally important features Braun et al. (2024); Makelov et al. (2024). To address this, Braun et al. (2024) proposed end-to-end (e2e) SAEs, which are trained to minimize the KL divergence between the output distributions of the original model and the model with SAE-generated activations inserted, instead of minimizing the reconstruction loss. This approach ensures that e2e SAEs are optimized to identify features that influence the model’s predictions, rather than focusing on features that only accurately reconstruct activations.
@@ 6 A Beginner’s Roadmap to MI
A key motivation for this survey is to provide a friendly guide for researchers and developers interested in MI to quickly pick up the field. To this end, we provide a beginner’s roadmap in Figure 9, 13, and 15, where we categorize MI research into three categories – feature study, circuit study, and the study of universality, corresponding to the three objects of studies (Section 4.1). Each category is further divided into sub-categories with their corresponding MI workflow and associated techniques.
@@ 6.1 Feature Study
A targeted feature study aims to determine whether a pre-defined feature is encoded in the representation of an LM. These pre-defined features are typically intuitive guesses or hypotheses made by humans, and we perform a targeted feature study when we want to verify the presence or absence of these features. Therefore, a general research question in a targeted feature study is, “Does this LM representation (or a subset of its neurons) encode the feature FF?”
@@ 6.1.1 General Workflow for Targeted Feature Study
A targeted feature study begins with a hypothesis about the presence or absence of a pre-defined feature FF in an LM representation. For example, suppose we are interested in knowing whether an LM activation encodes the “is_python_code” feature in its activations when the input contains a Python code snippet. Then, our hypothesis would be: “Does this LM activation (e.g., $h_{i}^{1}$) encode the ‘is_python_code’ feature?” We can refine our hypothesis to localize the feature encoding to a more granular level: “Does this neuron in this activation (e.g., Neuron 1 in $h_{i}^{1}$) encode ‘is_python_code’ feature?”. Note that we can iteratively pose this question to all the LM activations when we want to identify a complete set of activations that encode the targeted feature.
@@
To verify whether or not a pre-defined feature exists in a given LM representation, there are typically two workflows to consider. The first workflow directly probes the LM representation and inquires the presence of the targeted feature $F$, whereas the second workflow follows a similar procedure as the open-ended feature study, which first identifies all features encoded in the representation and then confirms whether the feature $F$ is among them. Below, we mainly present the first workflow. Details of the second workflow will be discussed in Section 6.1.2. For the first workflow, one option is to use the Probing method (Section 5.4.1) to verify the hypothesis, where a probe determining the existence of the targeted feature $F$ will be trained. Probing is a lightweight and easy-to-implement technique with a substantial body of literature discussing its strengths and limitations. PatchScope Ghandeharioun et al. (2024) is another technique that can also be used for verifying the hypothesis. This method is especially useful to verify whether certain attributes associated with a given subject (e.g., attributes associated with the “largest city of” from the representation of “United States”) are encoded in the LM representation. Unlike probing, these methods require no training. However, PatchScope relies on humans to craft an expressive inspection prompt capable of eliciting the desired feature, which can be particularly challenging for abstract features. Besides PatchScope, other vocabulary projection methods can also be used to check hypotheses about features that are expressible in the vocabulary space. Beyond these approaches, DII Geiger et al. (2024) offers a more fine-grained method for hypothesis verification. DII operates at the subspace level, enabling interventions when a hypothesis posits that a given causal variable is encoded not in the full activation but in a distinct subspace of the activation. Concretely, DII applies a change-of-basis transformation to rotate the representation, intervenes on a $k$-dimensional subspace corresponding to the high-level causal variable, and then maps the representation back to the original basis. Complementary to DII, DAS Geiger et al. (2024) is a supervised approach that automatically discovers both the change-of-basis transformation and the $k$-dimensional subspace best aligned with a given causal variable. In short, DII enables fine-grained causal interventions, while DAS identifies the appropriate subspaces in which those interventions should be applied. The hypothesis can be validated based on the intervention effect of DII or whether DAS can identify an appropriate subspace for the targeted causal variable. However, both DII and DAS require access to training data composed of clean-counterfactual pairs. Therefore, the choice of these techniques depends on the type of feature being investigated as well as the available computational and data resources. We also summarize the advantages and limitations of the two techniques in Table 7.
Techniques Advantages Limitations
Targeted Feature Study
Probing Established methodology with a large body of literature; useful for identifying a wide range of feature types (e.g., syntactic or semantic features). Require training data; correlational only (not causal evidence); confirms only feature presence (not usage).
PatchScope No additional training; lightweight. Require well-crafted inspection prompts; unsuitable for discovering abstract features (e.g., syntactic features).
Vocabulary Projection Methods No additional training; lightweight. Only decode features that can be represented in vocabulary space; early-layer projections can be unreliable.
Distributed Alignment Search (DAS) Discovers the specific $k$-dimensional subspace associated with a causal variable from a given representation. Dataset with clean-counterfactual pairs required; computationally expensive.
Open-ended Feature Discovery Can uncover unexpected related features. Computationally expensive; human-in-the-loop labeling.
Open-Ended Feature Study — Observe (Step 1)
Neuron Activation Visualization Intuitive; no additional training. Neurons encode polysemantic features that are challenging to interpret.
Sparse Autoencoders (SAE) Project activations to sparse activation with more monosemantic features, improving interpretability of activations. Require SAE training; high compute cost; sparsity–reconstruction trade-offs.
Vocabulary Projection Methods Provide important signals on how a given activation or feature impacts LM output. Only decode features that can be represented in vocabulary space; early-layer projections can be unreliable without translators.
Open-Ended Feature Study — Explain (Step 2)
Human Gold standard in current practice. Expensive and time-consuming; subjective; polysemantic features are hard to label consistently across annotators.
Automatic LM Reduce human effort; scalable first-pass explanations and scoring; consistent formatting. Faithfulness and reliability of explanations must be validated; risk of model bias or hallucination; may bias human reviewers.
Human + Automatic LM Efficient division of labor: LMs propose/cluster, humans verify/refine; improve efficiency while maintaining quality Still need human oversight; potential anchoring on LM suggestions.
Table 7: Comparison of techniques across Targeted and Open-ended feature study stages (Observe & Explain).
@@ 6.1.2 General Workflow for Open-ended Feature Study
In this step, we employ various MI techniques to gather various information about the activation to guide open-ended feature discovery. The gathered information can be referred to as “intermediate explanations” because they require further interpretation during the Explain step to discover the exact features. Techniques such as vocabulary projection methods (Section 5.1), neuron activation visualization (Section 5.4.2), and SAEs (Section 5.3) are commonly used techniques for generating intermediate explanations. We provide the advantages and limitations of these techniques in Table 7.
@@
Neuron activation visualization (Sec 5.4.2) is another technique that involves developing an interface to visually highlight the text or tokens that elicit information about a specific neuron. The idea here is that, for example, if a neuron consistently activates for text written in French, one can then label the neuron as encoding a French text detection feature. It is important to note that neuron activation visualization is based on an important assumption that features are represented by neurons in a one-to-one correspondence. However, subsequent studies Elhage et al. (2022a); Marks et al. (2024b) have found polysemantic neurons, i.e., neurons that activate for multiple unrelated features, rendering the assumption of one-to-one correspondence to be false. The discovery of polysemantic neurons also complicates the Explain step, since polysemantic neurons can easily confuse the explainer when they activate for multiple unrelated features simultaneously. To address this challenge, subsequent studies have proposed training SAEs (Section 5.3) that will transform the original activation into a higher-dimensional sparse activation. The goal of this projection is to generate a new representation that faithfully retains all the features from the original activation but encodes them in a way that each neuron corresponds to a single feature, resulting in a representation composed of more monosemantic neurons. The neuron activation visualization can then be applied to the sparse activation, rather than the original activation, for generating more interpretable intermediate explanations. However, training SAEs for specific LMs requires non-trivial computation, and SAEs, as we discussed in Section 5.3, face weaknesses such as the sparsity-reconstruction trade-off.
Although each of these techniques can be used individually to generate the intermediate explanation in the Observe step, in practice, they are often combined to provide a more complete explanation of the LM activation. Specifically, the intermediate explanations from various MI techniques are generated and presented in an interface for the human evaluators to annotate the feature description in the Explain step. In Section 6.1.3, we will discuss three examples of interfaces from prior work.